# Dimitrios M. Thilikos ${}^{,}$

### v3

@@ 5.1 A local structure theorem for excluding Dyck-grids
Let $\mathbb{S}$ be a non-empty finite and closed set of surfaces. Let $g_{1}$ be the Euler-genus of the orientable surface in $\mathsf{sobs}(\mathbb{S})$ and $g_{2}$ be the Euler-genus of the non-orientable surface in $\mathsf{sobs}(\mathbb{S})$ if such a surface exists, otherwise set $g_{2}\coloneqq 0.$
There exists a function $f\colon\mathbb{N}\to\mathbb{N}$ such that for every $k\in\mathbb{N},$ every graph $G$ which excludes the orientable Dyck-grid of order $k$ and Euler-genus $g_{1}$ and, if $g_{2}\neq 0,$ the non-orientable Dyck-grid of order $k$ and Euler-genus $g_{2}$ satisfies ${\mathbb{S}}\mbox{-}\mathsf{hw}(G)\leq f(k).$
In particular, $G$ has a tree-decomposition $(T,\beta)$ of adhesion less than $f(k)$ such that for every $d\in V(T),$ either $|\beta(d)|\leq f(k),$ or there exist a set $A_{d}\subseteq\beta(d)$ of size at most $f(k)$ and a surface $\Sigma\in\mathbb{S}$ such that $\mathsf{Torso}_{G}(d)-A_{d}$ has a $\Sigma$-decomposition $\delta$ of width at most $f(k)$ and breadth at most $f(k).$
[…]
  • i.
    for every neighbor $d^{\prime}$ of $d$ in $T$ we have that $|(\beta(d)\cap\beta(d^{\prime}))\setminus X_{d}|\leq 3$ and
  • ii.
    $\mathsf{Torso}_{G}(d)-X_{d}$ is a minor of $G.$

### v4

@@ 7 From local to global structure theorems
Let $k,$ $\mathsf{h}_{1},$ $\mathsf{h}_{2},$ $\mathsf{c}_{2},$ and $r$ be non-negative integers with $\mathsf{c}_{2}\in[0,2]$ and $0\leq\mathsf{h}_{2}<\mathsf{h}_{1}.$ Let $g_{1}\coloneqq 2\mathsf{h}_{1},$ $g_{2}\coloneqq 2\mathsf{h}_{2}+\mathsf{c}_{2}$ if $\mathsf{c}_{2}\neq 0$ and $g_{2}\coloneqq 0$ otherwise. Moreover, let $g\coloneqq\max\{g_{1},g_{2}\}$ and fix $t\coloneqq(162^{2g}+4)k.$
Then for every graph $G$ and every integer $r\geq 3$ one of the following is true.
  1. 1.
    $G$ contains a (a) $\mathscr{D}^{(\mathsf{h}_{1},0)}_{k}$ or a (b) $\mathscr{D}^{(\mathsf{h}_{2},\mathsf{c}_{2})}_{k}$-expansion $D$ where case (b) is allowed as an outcome if and only if $g_{2}\neq 0,$ or
  2. 2.
    $G$ has a tree decomposition $(T,\beta)$ of adhesion less than $\mathbf{a}(g_{1},g_{2},k,r)$ such that for every $d\in V(T),$ either $|\beta(d)|\leq 4\mathbf{a}(g_{1},g_{2},k,r),$ or there exists a set $A_{d}\subseteq\beta(d)$ of size at most $4\mathbf{a}(g_{1},g_{2},k,r)$ and integers $\mathsf{h}_{d},\mathsf{c}_{d},$ $\mathsf{c}_{d}\in[0,2],$ together with a surface $\Sigma=\Sigma^{(\mathsf{h}_{d},\mathsf{c}_{d})}$ where neither $\mathscr{D}^{(\mathsf{h}_{1},0)}_{i}$ nor, in case $\mathsf{c}_{2}\neq 0,$ $\mathscr{D}^{(\mathsf{h}_{2},\mathsf{c}_{2})}_{i}$ embeds for all $i\in\mathbb{N},$ such that for the torso $G_{d}$ of $G$ at $d,$ the graph $G_{d}-A_{d}$ has a $\Sigma$-decomposition $\delta$ of width at most $\mathbf{a}(g_{1},g_{2},k,r)$ and breadth at most $\mathbf{v}_{\mathsf{b},g_{1},g_{2}}$ such that
    1. (a)
      there exists an $r$-wall $W_{d}$ which is flat in $\delta,$
    2. (b)
      there is no vertex of $G_{d}-A_{d}$ which is drawn in the interior of a non-vortex cell of $\delta,$
    3. (c)
      $G_{d}-A_{d}$ is a minor of $G,$ and
    4. (d)
      for every neighbour $d^{\prime}$ of $d$ in $T$ it holds that $|\big{(}\beta(d)\cap\beta(d^{\prime})\big{)}\setminus A_{d}|\leq 3.$
Moreover, there exists a function $g\colon\mathbb{N}^{4}\to\mathbb{N}$ and an algorithm that finds one of these outcomes in time $g(k,\mathsf{h}_{1},r)|V(G)|^{4}\log(|V(G)|))$ where $g(k,\mathsf{h}_{1},r)=2^{2^{2^{\mathcal{O}(\mathsf{h}_{1})}\text{$\mathsf{poly}$}(k+r)}}.$